OpenStudy (anonymous):

12x3 + 8x2 + 4x 2. 12x2 - 7x - 10 3. 12x2 + 50xy + 8y2 4. 2x3 - 5x2y + xy2 + 2y3 5. 2x3 - 4x2y + xy2 - 2y3 a. Simplify (2x2 + y2)(x - 2y) = ? b. Simplify (6x + y)(2x + 8y )= ? c. Simplify (2x + y)(x2 -3xy + 2y 2 ) = ? d. Simplify (3x + 2)(4x - 5) = ? e. Simplify 4x(3x2+2x+1) = ?

OpenStudy (anonymous):

1. 12x3 + 8x2 + 4x 2. 12x2 - 7x - 10 3. 12x2 + 50xy + 8y2 4. 2x3 - 5x2y + xy2 + 2y3 5. 2x3 - 4x2y + xy2 - 2y3 a. Simplify (2x2 + y2)(x - 2y) = ? b. Simplify (6x + y)(2x + 8y )= ? c. Simplify (2x + y)(x2 -3xy + 2y 2 ) = ? d. Simplify (3x + 2)(4x - 5) = ? e. Simplify 4x(3x2+2x+1) = ?

OpenStudy (anonymous):

matching help

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Still there? All you need to do is distribute the term outside the parenthesis to every term within the parenthesis. Do you know how to do that?

OpenStudy (anonymous):

so a goes to 5.

OpenStudy (anonymous):

b goes to 3.

OpenStudy (anonymous):

c goes to four.

OpenStudy (anonymous):

d goes to 2.

OpenStudy (anonymous):

And e goes to 1 obviously.